how can i start solving this problem it saids solve the linear system x-4y=24 3x+y=7
Answer by dabanfield(803)

1.) x-4y=24
2.) 3x+y=7
Let's rewrite 1.) as x = 4y + 24
Then we can substitute 4y+24 for x in equation 2.):
3*(4y+24) + y = 7
12y + 72 + y = 7
13y = -65
y = -65/13
y = -5
Substutite -5 for y and calculate
x = 4*-5 + 24
x = -20 + 24
x = 4